# Energy-Aware Routing to Large Reasoning Models

### v1

@@ Abstract
Large reasoning models (LRMs) have heterogeneous inference energy costs based on which model is used and how much it reasons. To reduce energy, it is important to choose the right LRM and operate it in the right way. As a result, the performance of systems that dispatch tasks to different individual LRMs depend on the balance between mean energy provisioning and stochastic fluctuations. The critical regime is the unique operating point at which neither auxiliary energy nor baseline energy is systematically wasted. Increasing baseline supply shifts the system toward persistent over-supply and baseline-energy waste, while reducing supply induces persistent reliance on auxiliary energy. Yet in this regime, performance remains volatility-limited and so a second-order characterization provides further insights that we develop. Here, performance is governed by how variability is absorbed across time, models, and execution choices. This perspective highlights variance-aware routing and dispatch as a principled design axis, and provides a theoretical basis for developing energy-aware model routing policies. Routing behavior is characterized when dispatch policies are based on training-compute and inference-compute scaling laws for LRMs.
@@ I Introduction
Large artificial intelligence (AI) models have become widely used in the past few years, and recently large reasoning models (LRMs) have become prominent [PlaatWVBVB2025, EllisMohr2026TheoryInferenceCompute]. Such models are often infeasible for organizations to host locally due to hardware constraints, leading to the growth of large data centers—so-called AI factories—with massively parallel hardware to host an ever-growing number of models of varying complexity. These data centers expend a significant amount of energy to process model requests.
@@
There is recent interest in powering data centers with renewable energy [AgarwalSNIBCSK2021]. However, sources such as solar and wind introduce significant variability of available energy, which rarely aligns with usage. To mitigate variability, Varaiya et al. considered various ways to optimize risk-limited dispatch of energy for numerous physical workloads [Varaiya2011RiskLimitedDispatchEnergy], but the basic formalism can be extended to informational workloads [Varshney2014RiskLimitedDispatchKnowledge]. Here we aim to draw on the specific properties of LRMs in AI factories, where there is a possibility of routing tasks to particular LRMs [ShnitzerOSSSSTY2024]. Notably, the use and energy requirements of AI models also fluctuates with time [Luccioni2023bloom]. Moreover, AI models with varying capabilities have different inference energy costs [JEGHAMAKEH2025, TRIPPPGN2024, HUSOMGSS2024]. Often larger AI models yield better performance due to training-compute scaling [NayakV2026], but they require more energy to run. Further, due to inference-compute scaling of LRMs, more time/energy of computation may provide higher-quality responses [NayakV2026, EllisMohr2026TheoryInferenceCompute]. As such, there are two dimensions to energy optimization: which LRM to route to and how long to run the chosen LRM.
@@
Previous work has studied reducing energy requirements for large-scale AI systems in several ways. The Clover system experimentally showed that routing to a mixture of low- and high-quality models can improve energy efficiency, while maintaining performance [LISGT2023]. EcoServe focused on GPU and CPU usage optimization, specifically by exploiting underutilized host CPUs and dynamically scaling GPUs and CPUs [LIHCFSG2025]. Moreover, there are many works that focus on optimizing the AI models directly. FrugalGPT proposed the prompt adaption, LLM approximation, and LLM cascade strategies [ChenZZ2024]. Model pruning and knowledge distillation techniques have also been used for model efficiency [SunLBK2024, MURALIDKARANSJC2024, HsiehLYNF2024], which in turn decrease energy consumed. Although these types of approaches improve different aspects of AI factory efficiency, they typically study model inference efficiency and power system considerations in isolation. Therefore, this does not explicitly address deployment settings in which renewable energy availability, inference cost heterogeneity, and deadline constraints must be jointly considered.
@@
In this work, we introduce a mathematically principled formulation of the energy-aware model routing problem. This has formal similarity to information-theoretic investigations of optimal packet scheduling in energy harvesting systems [YANGS2012, UlukusYESZGH2015], in which packets and harvested energy arrive randomly, and the goal is to minimize the time to send data packets. There are several second-order characterizations of energy-harvesting channels, characterizing channel dispersion essentially using Berry-Esseen forms of the central limit theorem [MolavianJaziY2015], but we study first-order and second-order characterizations for our problem directly using properties of Brownian motion. Also, in our setup, renewable energy can be augmented with non-renewable energy to meet the task constraints (i.e., time and accuracy). The routing policy assigns each task to a hosted model with the objective of minimizing auxiliary energy consumption, subject to the request’s constraints.
@@
In addition to our key results on the first- and second-order characterizations for energy-aware routing to LRMs, we also provide deep connections to training-compute and inference-compute scaling laws that are empirically well-established and for which there are nascent theoretical explanations [Bahri2024explainingscaling, NayakV2026, EllisMohr2026TheoryInferenceCompute, Wadell2025FMChem]. Basing task dispatch on these scaling laws not only reduces the need for an energy-heavy dispatcher running a large AI model itself, but also provides practical guidance for dispatch policies in deployed AI factories.
@@ II System Model and Problem Formulation
Figure 1: System diagram
Consider an LRM routing system with unlimited parallel processing capacity: any number of tasks may be processed concurrently by any AI model, and the only binding resource is energy. The index set of available LRMs is $\mathcal{M}$. Tasks arrive stochastically and enter a router buffer (queue). At decision times, the router may dispatch any subset of queued tasks into processing (service). Once dispatched, a task remains in processing for a policy-chosen thinking time and consumes energy throughout that interval; it is released (and stops using energy) at completion. Energy availability is also governed by stochastic processes $R$ for baseline energy and $G$ for auxiliary energy. See Fig. 1.
@@ II-A Tasks, requirements, and arrivals
$x\coloneqq(\theta,r)\in\mathcal{X},\qquad r\coloneqq(\lambda,\varepsilon),$
@@ II-B Router actions and task life-cycle
$(i(x),\tau(x))=\pi(\text{state at time }s(x)).$
@@ II-C Reliability model and oracle stopping
We adopt a best-case oracle stopping assumption: for any chosen $\tau$, the model runs until time $\tau$ and halts immediately when instructed. Thus, once the router dispatches a task and chooses $(i,\tau)$ satisfying (3), the task is guaranteed to stop (and therefore stop consuming energy) at completion time $s(x)+\tau$. This isolates routing/dispatch limits from early-stopping design.
@@ II-D System energetics
If task $x=(\theta,(\lambda,\varepsilon))$ is dispatched to model $M_{i}$ for thinking time $\tau$, let $e_{i}(x,u)\geq 0$ denote its instantaneous energy-consumption rate at elapsed service time $u\in[0,\tau]$ (measured from the dispatch instant). The total energy required to run $x$ on $M_{i}$ for duration $\tau$ is
$E_{i}(x,\tau)=\sum_{u=1}^{\tau}e_{i}(x,u).$ (4)
Let $\mathcal{Q}(t)$ denote the set of tasks that have arrived by time $t$ and have not yet been serviced. Under a policy $\pi$, each task $x$ has a dispatch time $s(x)\geq t_{0}(x)$ and (chosen at dispatch) an assigned model $i(x)\in\mathcal{M}$ and thinking time $\tau(x)\geq 0$, inducing a service interval $s(x),\,s(x)+\tau(x))$. Define the set of tasks in service at time $t$ as
$\mathcal{S}(t)\coloneqq\{x:t_{0}(x)\leq t,s(x)\leq t<s(x)+\tau(x)\}.$
@@
$C_{t}\coloneqq\sum_{x\in\mathcal{S}_{t}}e_{i(x)}\bigl(x,\,t-s(x)\bigr),$ (5)
@@ II-E Objective and simplified dynamics
A policy $\pi$ observes the arriving tasks and system state (including $\widetilde{B}_{t}$ and optionally harvest side-information) and makes online dispatch decisions, chooses allocations $(i(x),\tau(x))$ for dispatched tasks, and selects auxiliary energy usage $G_{t}$ when needed. The goal is to satisfy, for every task, the deadline constraint (1) and the tolerance constraint (3), while minimizing reliance on auxiliary energy. Thus, considering the deficit at time $T$:
$D_{T}\coloneqq\sum_{t=1}^{T-1}G_{t}$ (7)
@@
$\min_{\pi}\ \operatorname{\mathbb{E}}_{\pi}\left[D_{T}\right],$
@@
Fix a horizon $T\in\mathbb{N}$, and exogenous sequences $\{R_{t}\}_{t=0}^{T-1}$ and $\{C_{t}\}_{t=0}^{T-1}$ in $\mathbb{R}$. Let $t=0,1,\dots,T-1$ and the uncontrolled (possibly negative) battery trajectory $\{B_{t}\}_{t=0}^{T}$ be defined by
$B_{0}\geq 0,$ (9)
@@
$\widetilde{B}_{0}\coloneqq B_{0},$ (11)
$\widetilde{B}_{t+1}\coloneqq\widetilde{B}_{t}+R_{t}-C_{t}+G_{t},$ (12)
where $G_{t}\geq 0$ for all $t$. Consider the greedy (minimal) choice
@@
$D_{T}=\sum_{t=0}^{T-1}G_{t}=\Bigl(-\min_{0\leq t\leq T}B_{t}\Bigr)^{+}.$
Then, the asymptotic objective with constraints is:
$\overline{J}_{\pi}^{\star}$ $\coloneqq\min_{\pi}\ \limsup_{T\to\infty}\ \mathbb{E}_{\pi}\!\left[\frac{1}{T}\Bigl(-\min_{0\leq t\leq T}B_{t}\Bigr)^{+}\right]$ (14)
@@ III Performance Analysis
We now characterize the fundamental limits of the routing system under ergodicity assumptions.
@@ III-A Ergodic arrivals and harvesting
We assume empty initial battery; task arrivals $\{(K_{t},\{X_{t,k}\}_{k=1}^{K_{t}})\}_{t\geq 1}$, where a task is now denoted by random variable $X_{t,k}$; and energy harvests $\{R_{t}\}_{t\geq 1}$ are mutually independent i.i.d. sequences such that:
$K_{t}\sim\mathrm{Poisson}(\overline{K}),$
$\operatorname{\mathbb{E}}[R_{t}]=\overline{R},\quad\operatorname{Var}(R_{t})=\sigma_{R}^{2},$
$X_{t,k}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}f_{X},\quad\forall\,t,k,$
$B_{0}=0.$
@@
be a (measurable) minimizer, and define $\tau_{\mathrm{LB}}(X)\coloneqq\tau^{*}_{i_{\mathrm{LB}}(X)}(X)$ and $e_{\mathrm{LB}}(X,u)\coloneqq e_{i_{\mathrm{LB}}(X)}(X,u)$ for $u=1,\dots,\tau_{\mathrm{LB}}(X)$, so that the lower bound on processing task $X$ is
$E_{\mathrm{LB}}(X)=E_{i_{\mathrm{LB}}(X)}^{*}(X)=\sum_{u=1}^{\tau_{\mathrm{LB}}(X)}e_{\mathrm{LB}}(X,u).$ (19)
@@
$\overline{C}_{\mathrm{LB}}$ $\coloneqq\operatorname{\mathbb{E}}\!\left[\sum_{k=1}^{K_{t}}E_{\mathrm{LB}}\!\left(X_{t,k}\right)\right]$
$\ =\overline{K}\,\operatorname{\mathbb{E}}_{X\sim f_{X}}\!\left[E_{\mathrm{LB}}(X)\right],$
where the later equality comes under Assumption 1 and independence of $K_{t}$ and $\{X_{t,k}\}_{k=1}^{K_{t}}$. This provides a policy-independent lower bound on the long-run average energy consumption rate.
@@ III-B Myopic dispatcher scaling analysis
We now analyze a simple baseline that charges each arriving task’s per-task energy lower bound in its arrival slot. This does not exploit the option to slide service within latency windows; it is therefore a tractable reference model for variability and reserve scaling.
We consider two myopic-in-time baselines that both begin service immediately upon arrival, but differ in how energy is accounted: (i) lumped-at-arrival consumption and (ii) distributed-in-service consumption. Define the lumped myopic per-slot consumption:
$C_{t}^{\mathrm{my,lump}}\coloneqq\sum_{k=1}^{K_{t}}E_{\mathrm{LB}}(X_{t,k}).$
@@
$C_{t}^{\mathrm{my,dist}}\coloneqq\sum_{x\in\mathcal{S}(t)}e_{\mathrm{LB}}\bigl(x,\,t-s(x)+1\bigr).$

### v2

@@ Abstract
Large reasoning models (LRMs) have heterogeneous inference energy costs based on which model is used and how much it reasons. To reduce energy, it is important to choose the right LRM and operate it in the right way. As a result, the performance of systems that dispatch tasks to different individual LRMs depend on the balance between mean energy provisioning and stochastic fluctuations. The critical regime is the unique operating point at which neither auxiliary energy nor baseline energy is systematically wasted. Increasing baseline supply shifts the system toward persistent over-supply and baseline-energy waste, while reducing supply induces persistent reliance on auxiliary energy. Yet in this regime, performance remains volatility-limited and so a second-order characterization provides further insights that we develop. Here, performance is governed by how variability is absorbed across time, models, and execution choices. This perspective highlights variance-aware routing and dispatch as a principled design axis, and provides a theoretical basis for developing energy-aware model routing policies. We characterize routing behavior under a simple dispatch policy based on training- and inference-compute scaling laws for LRMs.
@@ I Introduction
Large artificial intelligence (AI) models have become widely used in the past few years, and recently large reasoning models (LRMs) have become prominent [24, 9]. Such models are often infeasible for organizations to host locally due to hardware constraints, leading to the growth of large data centers—so-called AI factories—with massively parallel hardware to host an ever-growing number of models of varying complexity. These data centers expend a significant amount of energy to process model requests.
@@
There has been recent interest in powering data centers with renewable energy [1]. However, sources such as solar and wind introduce significant variability in available energy, which rarely aligns with usage. To mitigate variability, Varaiya et al. considered various ways to optimize risk-limited dispatch of energy for numerous physical workloads [30], which can be extended to informational workloads [2]. Here we aim to draw on the specific properties of LRMs in AI factories, where there is a possibility of routing tasks to particular LRMs [25]. Notably, the use and energy requirements of AI models also fluctuates with time [20]. Moreover, AI models with varying capabilities have different inference energy costs [14, 28, 13]. Larger AI models tend to yield better performance due to training-compute scaling [23], but this typically comes with higher inference-time energy use. Further, due to inference-compute scaling of LRMs, more time/energy of computation may provide higher-quality responses [23, 9]. As such, there are two main dimensions of energy optimization: which LRM to route to and how long to run it for a given workload.
@@
Previous work has studied reducing energy requirements for large-scale AI systems in several ways. The Clover system experimentally showed that routing to a mixture of low- and high-quality models can improve energy efficiency, while maintaining performance [18]. EcoServe focused on GPU and CPU usage optimization, specifically by exploiting underutilized host CPUs and dynamically scaling GPUs and CPUs [19]. Moreover, there are many works that focus on optimizing AI models directly. FrugalGPT proposed the prompt adaption, LLM approximation, and LLM cascade strategies [6]. Model pruning and knowledge distillation techniques have also been used for model efficiency [26, 22, 12], which in turn decrease the energy consumed. Although these types of approaches improve different aspects of AI factory efficiency, they typically study model inference efficiency and power system considerations in isolation. Therefore, this does not explicitly address deployment settings in which renewable energy availability, inference cost heterogeneity, and deadline constraints must be jointly considered.
@@
In this work, we introduce a mathematically principled formulation of the energy-aware model routing problem. This has formal similarity to information-theoretic investigations of optimal packet scheduling in energy harvesting systems [32, 29], in which packets and harvested energy arrive randomly, and the goal is to minimize the time to send data packets. There are several second-order characterizations of energy-harvesting channels that characterize channel dispersion using Berry-Esseen forms of the central limit theorem [21], but we study first-order and second-order characterizations for our problem directly using properties of Brownian motion. Also, in our setup, renewable energy can be augmented with non-renewable energy to meet the task constraints (i.e., time and accuracy). The routing policy assigns each task to a hosted model with the objective of minimizing auxiliary energy consumption, subject to the request’s constraints.
@@
In addition to our key results on the first- and second-order characterizations of energy-aware routing to LRMs, we also provide deep connections to training-compute and inference-compute scaling laws that are empirically well-established and for which there are nascent theoretical explanations [3, 23, 9, 31]. Basing task dispatch on these scaling laws not only reduces the need for an energy-heavy dispatcher running a large AI model itself, but also provides practical guidance for dispatch policies in deployed AI factories.
@@ II System Model and Problem Formulation
Figure 1: System diagram of the LRM routing system and energy flow: incoming tasks join the pending set, the dispatcher selects a feasible model and runtime for each dispatched task, and the battery supplies energy for execution, drawing auxiliary energy when needed.
Consider an LRM routing system with unlimited parallel processing capacity: any number of tasks may be processed concurrently by any AI model, and the only binding resource is energy. The index set of available LRMs is $\mathcal{M}$. Tasks arrive stochastically and enter a router buffer (queue). At decision times, the router may dispatch any subset of queued tasks into processing (service). Any number of queued tasks may be dispatched to the same model $M_{i}\in\mathcal{M}$, even at the same time. Once dispatched, a task remains in processing for a policy-chosen thinking time and consumes energy throughout that interval; it is released (and stops using energy) at completion. Energy availability is also governed by stochastic processes $R$ for baseline energy and $G$ for auxiliary energy. See Fig. 1.
@@ II-A Tasks, requirements, and arrivals
$x\coloneqq(\theta,r)\in\mathcal{X},\qquad r\coloneqq(\lambda,\varepsilon),$ (1)
@@ II-B Router actions and task life-cycle
$(i(x),\tau(x))=\pi(\text{state at time }s(x)).$ (2)
@@ II-C Reliability model and oracle stopping
We adopt a best-case oracle stopping assumption: for any chosen $\tau$, the model runs until time $\tau$ and halts immediately when instructed. Thus, once the router dispatches a task and chooses $(i,\tau)$ satisfying (5), the task is guaranteed to stop (and therefore stop consuming energy) at completion time $s(x)+\tau$. This isolates routing/dispatch limits from early-stopping design. Termination can be implemented by enforcing a fixed compute budget while optionally triggering an earlier stop signal (for example by injecting a forced end-of-thinking token with reserved slack to finalize the response), with confidence- or verifier-based criteria as additional layers.
@@ II-D System energetics
If task $x=(\theta,(\lambda,\varepsilon))$ is dispatched to model $M_{i}$ for thinking time $\tau$, let $e_{i}(x,u)\geq 0$ denote its instantaneous energy-consumption rate at elapsed service time $u\in\{0,1,...,\tau-1\}$ (measured from the dispatch instant). The total energy required to run $x$ on $M_{i}$ for duration $\tau$
$E_{i}(x,\tau)=\sum_{u=0}^{\tau-1}e_{i}(x,u).$ (6)
Let $\mathcal{Q}(t)$ denote the set of tasks that have arrived by time $t$ and have not yet been dispatched. Under a policy $\pi$, each task $x$ has a dispatch time $s(x)\geq t_{0}(x)$ and (chosen at dispatch) an assigned model $i(x)\in\mathcal{M}$ and thinking time $\tau(x)\geq 0$, inducing a service interval $s(x),\,s(x)+\tau(x))$. Define the set of tasks in service at time $t$ as
$\mathcal{S}(t)\coloneqq\{x:t_{0}(x)\leq t,s(x)\leq t<s(x)+\tau(x)\}.$ (7)
@@
$C_{t}\coloneqq\sum_{x\in\mathcal{S}(t)}e_{i(x)}\bigl(x,\,t-s(x)\bigr),$ (8)
@@
Our goal here is to characterize system-level feasibility and cost structures under baseline variability, while making the dependence on task requirements and routing decisions explicit. We adopt abstractions that isolate the dispatcher layer from other mechanisms. While modern inference engines support token- and iteration-level scheduling and sophisticated memory management [16], we model each dispatched task as running continuously once started in order to focus on dispatcher-layer routing decisions without introducing additional engine-level coordination state. Accordingly, we abstract away mechanisms such as batching coordination across concurrent requests, queue reordering within the engine, and KV-cache management. Preemption may be incorporated by refining the time discretization and introducing micro-step scheduling (with cache and overhead modeling), which changes the set of feasible induced trajectories $\{C_{t}\}$ but not the form of the $\{R_{t}\}$$\{C_{t}\}$ analysis provided.
@@ II-E Objective and simplified dynamics
A policy $\pi$ observes the arriving tasks and system state (including $\widetilde{B}_{t}$ and optionally harvest side-information) and makes online dispatch decisions, chooses allocations $(i(x),\tau(x))$ for dispatched tasks, and selects auxiliary energy usage $G_{t}$ when needed. In addition to observing realized harvests $\{R_{t}\}$, the dispatcher may observe exogenous covariates that capture predictable spatio-temporal structure in renewable availability. The goal is to satisfy, for every task, the deadline constraint (3) and the tolerance constraint (5), while minimizing reliance on auxiliary energy. Thus, considering the deficit at time $T$:
$D_{T}\coloneqq\sum_{t=0}^{T-1}G_{t}$ (10)
@@
$\min_{\pi}\ \operatorname{\mathbb{E}}_{\pi}\left[D_{T}\right],$ (11)
@@
Fix a horizon $T\in\mathbb{N}$, and exogenous sequences $\{R_{t}\}_{t=0}^{T-1}$ and $\{C_{t}\}_{t=0}^{T-1}$ in $\mathbb{R}$. Let $t=0,1,\dots,T-1$, and let the uncontrolled (possibly negative) battery trajectory $\{B_{t}\}_{t=0}^{T}$ be defined by
@@
$\widetilde{B}_{t+1}\coloneqq\widetilde{B}_{t}+R_{t}-C_{t}+G_{t},\quad B_{0}\geq 0.$ (14)
where $G_{t}\geq 0$ for all $t$ with $\widetilde{B}_{0}=B_{0}$. Consider the greedy (minimal) choice
@@
$D_{T}=\sum_{t=0}^{T-1}G_{t}=\Bigl(-\min_{0\leq t\leq T}B_{t}\Bigr)^{+}.$ (16)
Thm. 1 is a pathwise identity for the harvested-energy sequence $\{R_{t}\}$ and the consumption sequence $\{C_{t}\}$; all routing-policy dependence enters through the induced consumption $C_{t}$ as defined in Eq. (8). Then, the asymptotic objective with constraints is:
$\overline{J}^{\star}$ $\coloneqq\min_{\pi}\ \limsup_{T\to\infty}\ \mathbb{E}_{\pi}\!\left[\frac{1}{T}\Bigl(-\min_{0\leq t\leq T}B_{t}\Bigr)^{+}\right]$ (17)
@@ III Performance Analysis
We now characterize the fundamental limits of the routing system under ergodicity assumptions to demonstrate cost factors under an explicit myopic policy realization.
@@ III-A Ergodic arrivals and harvesting
We assume an empty initial battery, i.e., $B_{0}=0$. Task arrivals $\{(K_{t},\{X_{t,k}\}_{k=1}^{K_{t}})\}_{t\geq 1}$ follow a Poisson process, where $K_{t}\sim\mathrm{Poisson}(\overline{K})$ and each task is represented by the random variable $X_{t,k}$. The task attributes $\{X_{t,k}\}$ are i.i.d. with distribution $f_{X}$. Energy harvests $\{R_{t}\}_{t\geq 1}$ form a mutually independent i.i.d. sequence with mean $\mathbb{E}[R_{t}]=\overline{R}$ and variance $\mathrm{Var}(R_{t})=\sigma_{R}^{2}$. Moreover, the sequences $\{K_{t}\}$, $\{X_{t,k}\}$, and $\{R_{t}\}$ are mutually independent.
Remark 1 (Interpretation of Assumption 1)
In practice, $R_{t}$ may be nonstationary or spatiotemporally correlated through exogenous covariates. Assumption 1 is adopted to enable a tractable second-order (drift–fluctuation) analysis. Accordingly, our i.i.d. model can be viewed as a baseline for the short-timescale supply variability that remains after accounting for predictable structure (e.g., diurnal/seasonal components or location-dependent means via forecasting, planning, or workload shifting). We leave explicit correlated and location-aware energy models to future work.
@@
be a (measurable) minimizer, and define $\tau_{\mathrm{LB}}(X)\coloneqq\tau^{*}_{i_{\mathrm{LB}}(X)}(X)$ and $e_{\mathrm{LB}}(X,u)\coloneqq e_{i_{\mathrm{LB}}(X)}(X,u)$ for $u=\{0,\dots,\tau_{\mathrm{LB}}(X)-1\}$, so that the lower bound on processing task $X$ is
$E_{\mathrm{LB}}(X)=E_{i_{\mathrm{LB}}(X)}^{*}(X)=\sum_{u=0}^{\tau_{\mathrm{LB}}(X)-1}e_{\mathrm{LB}}(X,u).$ (22)
@@
$\overline{C}_{\mathrm{LB}}\coloneqq\operatorname{\mathbb{E}}\Bigl[\sum_{k=1}^{K_{t}}E_{\mathrm{LB}}\!\left(X_{t,k}\right)\Bigr]=\overline{K}\,\operatorname{\mathbb{E}}_{X}\!\left[E_{\mathrm{LB}}(X)\right],$ (23)
where the latter equality comes under Assumption 1 and independence of $K_{t}$ and $\{X_{t,k}\}_{k=1}^{K_{t}}$. This provides a policy-independent lower bound on the long-run average energy consumption rate.
@@ III-B Myopic dispatcher scaling analysis
We now analyze a myopic baseline policy, $\pi_{\mathrm{my}}$, that dispatches each task in its arrival slot to the paired respective energy lower bound model. This does not exploit the option to slide service within latency windows; it is therefore a tractable reference model for variability and reserve scaling. We consider two myopic-in-time baselines that both begin service immediately upon arrival, but differ in how energy is accounted: (i) lumped-at-arrival consumption and (ii) distributed-in-service consumption. Define the lumped myopic per-slot consumption:
$C_{t}^{\mathrm{my,lump}}\coloneqq\sum_{k=1}^{K_{t}}E_{\mathrm{LB}}(X_{t,k}).$ (24)
@@
$C_{t}^{\mathrm{my,dist}}\coloneqq\sum_{x\in\mathcal{S}(t)}e_{\mathrm{LB}}\bigl(x,\,t-s(x)\bigr).$ (25)